Consider the following reaction:

Mg(s) + 2 HC1(aq)= MgC12(aq) + H2 (g)


The total pressure of gas collected over water is 785.0 mmHg and the temperature is 25.0C. What is the pressure of hydrogen gas formed in mmHg?


1
Expert's answer
2021-04-30T07:42:15-0400

"P_1=785mmHg"

"P_2=unknown"

"T_1=273k"

"T_2=298k"

From the ideal gas equation;(Gay lussacs law)

"P_2=P_1" "\times" "V\over V" "\times" "T\over T" "=P_1" "\times" "T\over T"

"=785mmHg\times" "298k\over 273k"

"=856.89mmHg"